The width of rectangular field is half its length. if its area is 512 sq.m find its perimeter.
Answer by TimothyLamb(4379)

w = L/2
L = 2w
Lw = 512
---
Lw = 512
2ww = 512
ww = 256
w = 16
L = 32
---
2L + 2w = 64 + 32
2L + 2w = 96
---
answer:
perimeter = 96 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php